Jeremy had 12 coins. 1/2 of the coins were quarters. 1/4 of the coins were dimes, and the rest were nickels. How much money does Jeremy have?
Answer by mananth(12270)

jermy has 12 coins
1/2 were quarters = 6
1/4 were dimes = 3
balance nickels = 3
6*25 + 3*10+3*5 = 195 cents
Ananth